# Rigidity results for initial data sets satisfying the dominant energy condition

### v3

@@ 4.4. Proof of Theorem I in the simply connected case
for each $i=1,\ldots,n-1$. By Proposition 4.1, we know that $q(e_{i},e_{j})=0$ for $i,j=1,\ldots,n-1$. This gives
00 0 $=\sum_{j=1}^{n-1}h(e_{i},e_{j})\,\langle T\wedge e_{j},\nu\wedge e_{k}\rangle_{\mathbb{R}^{n,1}}$
$=\langle b_{0}\wedge dB(e_{i}),\nu\wedge e_{k}\rangle_{\mathbb{R}^{n,1}}$
$=\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\,\langle dB(e_{i}),e_{k}\rangle_{\mathbb{R}^{n,1}}.$
Hence, if $\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\neq 0$ at the point $p$, then $dB=0$ at the point $p$. This implies $dN=0$ at the point $p$. Using Proposition 4.1, it follows that $h=0$ at the point $p$. ∎
@@
The composition of the isometric immersion $F\colon(M,g)\to\mathbb{R}^{n,1}$ with the projection $\pi\colon\mathbb{R}^{n,1}\to\mathbb{R}^{n}$, $(t,x)\mapsto x$, is a diffeomorphism onto a compact convex domain $K\subset\mathbb{R}^{n}$ with smooth boundary. In particular, $F$ is an isometric embedding onto the graph of a smooth real valued function defined on $K$.
@@
for $X,Y\in TM$. In other words, $g$ is the pull-back of the Euclidean metric under the immersion $\pi\circ F\colon M\to\mathbb{R}^{n}$. In particular, the metric $g_{0}$ is flat. It follows from Proposition 4.29 that the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ coincides with the second fundamental form of $\partial M$ with respect to the metric $g$. Using Proposition 4.1, we conclude that the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ is positive semi-definite at each point on $\partial M$. Moreover, since $N$ has non-zero degree, we can find a point on $\partial M$ where the Jacobian determinant of $N$ is non-zero, and at that point the second fundamental of $\partial M$ with respect to $g_{0}$ is positive definite. Theorem A.1 now implies that there exists an isometry of $(M,g_{0})$ onto a compact convex domain in $\mathbb{R}^{n}$ with smooth boundary. The isometric immersion $\pi\circ F\colon(M,g_{0})\to\mathbb{R}^{n}$ coincides with this isometry up to a rigid motion of $\mathbb{R}^{n}$. This implies the assertion. ∎
@@ 4.5. Proof of Theorem I in the general case
In this final subsection, we treat the case when $M$ is not simply connected. In this case, the isometric immersion $F:M\to\mathbb{R}^{n,1}$ is only locally defined, but its differential its globally defined. This is sufficient to define the bundle $\mathcal{W}$, the connection $\nabla^{\mathcal{W}}$, and the map $\gamma^{\mathcal{W}}:\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$. The arguments in Subsection 4.4 now imply that there exists a vector $b_{0}\in\mathbb{R}^{n,1}$ such that $\langle b_{0},b_{0}\rangle_{\mathbb{R}^{n,1}}=-1$ and $\langle b_{0},dF(X)\rangle_{\mathbb{R}^{n,1}}=0$ for every tangent vector $X\in T\Sigma$.
We now define a Riemannian metric $g_{0}$ on $M$ by
@@
for $X,Y\in TM$. Note that $g_{0}$ is a globally defined metric on $M$. As in Section 4.4, we see that $g_{0}$ is flat. Moreover, the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ coincides with the second fundamental form of $\partial M$ with respect to the metric $g$. In particular, the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ is positive semi-definite at each point on $\partial M$, and we can find a point on $\partial M$ where the second fundamental of $\partial M$ with respect to $g_{0}$ is positive definite. Using Theorem A.1, it follows that $M$ is diffeomorphic to a ball. This contradicts the assumption that $M$ is not simply connected.
5. Approximating a compact, convex polytope by smooth domains
Throughout this section, we fix an integer $n\geq 2$. We further assume that $\Omega\subset\mathbb{R}^{n}$ is a compact, convex polytope with non-empty interior. We write $\Omega=\bigcap_{i\in I}\{u_{i}\leq 0\}$, where $I$ is a finite set and the $u_{i}$ are linear functions on $\mathbb{R}^{n}$. After eliminating redundant inequalities, we may assume that the following condition is satisfied.
@@ 5. Approximating a compact, convex polytope by smooth domains
For each $i_{0}\in I$, the set
$\{u_{i_{0}}>0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}\leq 0\}$
is non-empty.
Let $g$ be a Riemannian metric which is defined on an open set containing $\Omega$, and let $q$ be a symmetric $(0,2)$-tensor. For each $i\in I$, $\nabla u_{i}$ will denote the gradient of $u_{i}$ with respect to the metric $g$, $D^{2}u_{i}$ will denote the Hessian of $u_{i}$ with respect to the metric $g$, $|\nabla u_{i}|_{g}$ will denote the norm of the gradient of $u_{i}$ with respect to the metric $g$, and $\nu_{i}=\frac{\nabla u_{i}}{|\nabla u_{i}|_{g}}$ will denote the unit normal vector field, with respect to the metric $g$, to the level sets of $u_{i}$. For each $i\in I$, we denote by $N_{i}\in S^{n-1}$ the outward-pointing unit normal vector to the halfspace $\{u_{i}\leq 0\}$ with respect to the Euclidean metric.
@@
If $\lambda$ is sufficiently large, then $\Omega_{\lambda}$ is a compact, convex domain in $\mathbb{R}^{n}$ with smooth boundary $\Sigma_{\lambda}=\partial\Omega_{\lambda}$. The sets $\Omega_{\lambda}$ form an increasing family of sets. Moreover,
@@
In the remainder of this section, $|\cdot|_{g}$ will denote the norm taken with respect to the metric $g$, and ||$\cdot$|| ⋅ | will denote the norm taken with respect to the Euclidean metric. The outward-pointing unit normal vector to the domain $\Omega_{\lambda}$ with respect to the metric $g$ is given by
@@
The map $N\colon\Sigma_{\lambda}\to S^{n-1}$ is homotopic to the Gauss map of $\Sigma_{\lambda}$ with respect to the Euclidean metric. A homotopy is induced by deforming $g$ to the Euclidean metric. In particular, the map $N$ has degree $1$.
Proposition 5.1 (Prop. 3.9 in [Brendle]).
@@
We define a function $W_{\lambda}\colon\Sigma_{\lambda}\to\mathbb{R}$ by
$W_{\lambda}=V_{\lambda}-|\mathrm{tr}(q)-q(\nu,\nu)|$
@@
Suppose that the Matching Angle Hypothesis is satisfied. Then $\sup_{\Sigma_{\lambda}}\max\{-W_{\lambda},0\}\leq o(\lambda)$ as $\lambda\to\infty$.
@@
It follows from Proposition 3.10 in [Brendle] that $\sup_{\Sigma_{\lambda}}\max\{-V_{\lambda},0\}\leq o(\lambda)$ as $\lambda\to\infty$. Since $W_{\lambda}-V_{\lambda}$ is uniformly bounded, the assertion follows. ∎
In the remainder of this section, we will estimate the $L^{\sigma}$-norm $\max\{-W_{\lambda},0\}$ on $\Sigma_{\lambda}\cap B_{r}(p)$, where $\sigma\in[1,\frac{3}{2})$ is a fixed exponent and $B_{r}(p)$ denotes a Euclidean ball of radius $r$. We begin by recalling a basic fact about the area of convex hypersurfaces in $\mathbb{R}^{n}$.
@@
Consider three pairwise distinct elements $i_{1},i_{2},i_{3}\in I$. We denote by $G_{\lambda}^{(i_{1},i_{2},i_{3})}$ the set of all points $x\in\Sigma_{\lambda}$ with the property that $u_{i_{1}}(x)\geq u_{i_{2}}(x)\geq u_{i_{3}}(x)$ and $u_{i_{3}}(x)\geq u_{i}(x)$ for $i\in I\setminus\{i_{1},i_{2},i_{3}\}$.
@@
For each $i\in I$, we assume that the hypersurface $\{u_{i}=0\}$ satisfies the condition $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$. Let us fix an exponent $\sigma\in[1,\frac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$. If $\lambda r$ is sufficiently large, then
@@
We adapt the proof of Lemma 3.13 in [Brendle]. Let us consider an arbitrary point $x\in G_{\lambda}^{(i_{1},i_{2},i_{3})}$ with $u_{i_{2}}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$. By definition of $G_{\lambda}^{(i_{1},i_{2},i_{3})}$, it follows that $u_{i}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$ for all $i\in I\setminus\{i_{1}\}$. Using the identity $\sum_{i\in I}e^{\lambda u_{i}(x)}=1$, we obtain $e^{\lambda u_{i_{1}}(x)}\geq 1-C\,e^{-(\lambda r)^{\frac{1}{8}}}$, hence $u_{i_{1}}(x)\geq-C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$. Arguing as in [Brendle], we obtain
@@
at the point $x$. Since $u_{i_{1}}(x)\geq-C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$ and $u_{i}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$ for all $i\in I\setminus\{i_{1}\}$, we can find a point $y\in\Omega$ such that $u_{i_{1}}(y)=0$ and $d(x,y)\leq C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$. By assumption, $H-|\mathrm{tr}(q)-q(\nu_{i_{1}},\nu_{i_{1}})|\geq 0$ at the point $y$. This implies
@@
for each point $x\in G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}$. On the other hand, the intersection $\Sigma_{\lambda}\cap B_{r}(p)$ has area at most $Cr^{n-1}$. Consequently,
@@
Assume that the Matching Angle Hypothesis holds. Let us fix an exponent $\sigma\in[1,\frac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$. If $\lambda r$ is sufficiently large, then
$\bigg{(}r^{\sigma+1-n}\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\geq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}\cap\{u_{i_{3}}\leq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}$
$\leq C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}$
@@
The corresponding estimate for $V_{\lambda}$ follows from Lemma 3.14 in [Brendle]. We next observe that $|W_{\lambda}-V_{\lambda}|\leq C\leq C\lambda^{\frac{1}{8}}r^{-\frac{7}{8}}$. Moreover, it is shown in [Brendle] that the set $G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\geq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}\cap\{u_{i_{3}}\leq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}$ has area at most $C\,(\lambda r)^{-\frac{7}{8}}\,r^{n-1}$. From this, the assertion follows. ∎
@@
Let us fix an exponent $\sigma\in[1,\frac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$. If $\lambda r$ is sufficiently large, then
@@
The corresponding estimate for $V_{\lambda}$ follows from Lemma 3.15 in [Brendle]. We next observe that $|W_{\lambda}-V_{\lambda}|\leq C\leq C\lambda$. Moreover, it is shown in [Brendle] that the set $G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{3}}\geq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}\cap B_{r}(p)$ has area at most $C\,(\lambda r)^{-\frac{3}{2}}\,r^{n-1}$. From this, the assertion follows. ∎
@@
For each $i\in I$, we assume that the hypersurface $\{u_{i}=0\}$ satisfies the condition $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$. Moreover, we assume that the Matching Angle Hypothesis is satisfied. Let us fix an exponent $\sigma\in[1,\frac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$.
If $\lambda r$ is sufficiently large, then
$\bigg{(}r^{\sigma+1-n}\int_{\Sigma_{\lambda}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}$
$\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}+C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}+C\,(\lambda r)^{1-\frac{3}{2\sigma}}.$
@@
$\bigg{(}r^{\sigma+1-n}\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}$
$\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}+C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}+C\,(\lambda r)^{1-\frac{3}{2\sigma}}$
@@
For each $i\in I$, we assume that the hypersurface $\{u_{i}=0\}$ satisfies the condition $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$. Moreover, we assume that the Matching Angle Hypothesis is satisfied. Let us fix an exponent $\sigma\in[1,\frac{3}{2})$.
Then
$\sup_{p\in\mathbb{R}^{n}}\sup_{r\leq 1}\bigg{(}r^{\sigma+1-n}\int_{\Sigma_{\lambda}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}\to 0$
as $\lambda\to\infty$.
@@
$(\delta_{l}\lambda_{l})^{-1}\sup_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}\to 0$
as $l\to\infty$. This implies

### v4

@@ 4.4. Proof of Theorem I in the simply connected case
for each $i=1,\ldots,n-1$. Proposition 4.1 implies that $q(e_{i},e_{j})=0$ for $i,j=1,\ldots,n-1$. This gives
$0$ $=\sum_{j=1}^{n-1}h(e_{i},e_{j})\,\langle T\wedge e_{j},e_{k}\wedge\nu\rangle_{\mathbb{R}^{n,1}}+\sum_{j=1}^{n-1}q(e_{i},e_{j})\,\langle e_{j}\wedge\nu,e_{k}\wedge\nu\rangle_{\mathbb{R}^{n,1}}$
$=\langle b_{0}\wedge dB(e_{i}),e_{k}\wedge\nu\rangle_{\mathbb{R}^{n,1}}$
$=-\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\,\langle dB(e_{i}),e_{k}\rangle_{\mathbb{R}^{n,1}}$
for $i,k=1,\ldots,n-1$. By assumption, $\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\neq 0$ at the point $p$. This implies $dB=0$ at the point $p$. Consequently, $dN=0$ at the point $p$. Using Proposition 4.1, we conclude that $h=0$ at the point $p$. Since $q^{\text{\rm tan}}=0$ by Proposition 4.1, it follows that the second fundamental form of $\Sigma$ in $\mathbb{R}^{n,1}$ vanishes at the point $p$. ∎
@@
Let $\pi$ denote the projection from $\mathbb{R}^{n,1}$ to
$\mathbb{R}^{n}=\{b\in\mathbb{R}^{n,1}\colon\langle b_{0},b\rangle_{\mathbb{R}^{n,1}}=0\}.$
The composition of the isometric immersion
$F\colon(M,g)\to\mathbb{R}^{n,1}$
with the projection $\pi\colon\mathbb{R}^{n,1}\to\mathbb{R}^{n}$ is a diffeomorphism from $M$ onto a compact convex domain $K\subset\mathbb{R}^{n}$ with smooth boundary. In particular, $F$ is an isometric embedding onto the graph of a smooth real valued function defined on $K$.
@@
for $X,Y\in TM$. In other words, $g_{0}$ is the pull-back of the Euclidean metric under the immersion $\pi\circ F\colon M\to\mathbb{R}^{n}$. In particular, the metric $g_{0}$ is flat. It follows from Proposition 4.31 that the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ coincides with the second fundamental form of $\partial M$ with respect to the metric $g$. Using Proposition 4.1, we conclude that the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ is weakly positive definite at each point on $\partial M$. Moreover, since $N$ has non-zero degree, we can find a point on $\partial M$ where the Jacobian determinant of $N$ is non-zero, and at that point the second fundamental of $\partial M$ with respect to $g_{0}$ is positive definite by Proposition 4.1. Theorem A.1 now implies that there exists an isometry of $(M,g_{0})$ onto a compact convex domain in $\mathbb{R}^{n}$ with smooth boundary. The isometric immersion $\pi\circ F\colon(M,g_{0})\to\mathbb{R}^{n}$ coincides with this isometry up to a rigid motion of $\mathbb{R}^{n}$. This implies the assertion. ∎
@@ 4.5. Proof of Theorem I in the general case
In this final subsection, we treat the case when $M$ is not simply connected. In this case, the isometric immersion $F\colon M\to\mathbb{R}^{n,1}$ is only locally defined, but its differential is globally defined. This is sufficient to define the bundle $\mathcal{W}$, the connection $\nabla^{\mathcal{W}}$, and the map $\gamma^{\mathcal{W}}\colon\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$. The arguments in Section 4.4 now imply that there exists a vector $b_{0}\in\mathbb{R}^{n,1}$ such that
$\langle b_{0},b_{0}\rangle_{\mathbb{R}^{n,1}}=-1\quad\text{ and }\quad\langle b_{0},dF(X)\rangle_{\mathbb{R}^{n,1}}=0$
for every tangent vector $X\in T\Sigma$.
As above, we define a Riemannian metric $g_{0}$ on $M$ by
@@
for $X,Y\in TM$. Note that $g_{0}$ is a globally defined metric on $M$. As in Section 4.4, we see that $g_{0}$ is flat. Moreover, the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ coincides with the second fundamental form of $\partial M$ with respect to the metric $g$. In particular, the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ is weakly positive definite at each point on $\partial M$, and we can find a point on $\partial M$ where the second fundamental of $\partial M$ with respect to $g_{0}$ is positive definite. Using Theorem A.1, it follows that $M$ is diffeomorphic to a ball. This contradicts the assumption that $M$ is not simply connected.
5. Approximating a compact convex polytope by smooth domains
Throughout this section, we fix an integer $n\geq 2$. We further assume that $\Omega\subset\mathbb{R}^{n}$ is a compact convex polytope with nonempty interior. We write $\Omega=\bigcap_{i\in I}\{u_{i}\leq 0\}$, where $I$ is a finite set and $u_{i}$, $i\in I$, is a collection of linear functions on $\mathbb{R}^{n}$. After eliminating redundant inequalities, we may assume that the following condition is satisfied.
@@ 5. Approximating a compact convex polytope by smooth domains
For each $i_{0}\in I$, the following set is nonempty:
$\{u_{i_{0}}>0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}\leq 0\}.$
Let $g$ be a Riemannian metric which is defined on an open set containing $\Omega$, and let $q$ be a symmetric $(0,2)$-tensor. For each $i\in I$, $\nabla u_{i}$ will denote the gradient of $u_{i}$ with respect to the metric $g$, $D^{2}u_{i}$ will denote the Hessian of $u_{i}$ with respect to the metric $g$, $|\nabla u_{i}|_{g}$ will denote the norm of the gradient of $u_{i}$ with respect to the metric $g$, and $\nu_{i}={\nabla u_{i}}/{|\nabla u_{i}|_{g}}$ will denote the unit normal vector field, with respect to the metric $g$, to the level sets of $u_{i}$. For each $i\in I$, we denote by $N_{i}\in S^{n-1}$ the outward-pointing unit normal vector to the halfspace $\{u_{i}\leq 0\}$ with respect to the Euclidean metric.
@@
If $\lambda$ is sufficiently large, then $\Omega_{\lambda}$ is a compact convex domain in $\mathbb{R}^{n}$ with smooth boundary $\Sigma_{\lambda}=\partial\Omega_{\lambda}$. The sets $\Omega_{\lambda}$ form an increasing family of sets. Moreover,
@@
In the remainder of this section, $|\cdot|_{g}$ will denote the norm taken with respect to the metric $g$, and $|\cdot|$ will denote the norm taken with respect to the Euclidean metric. The outward-pointing unit normal vector to the domain $\Omega_{\lambda}$ with respect to the metric $g$ is given by
@@
The map $N\colon\Sigma_{\lambda}\to S^{n-1}$ is homotopic to the Gauss map of $\Sigma_{\lambda}$ with respect to the Euclidean metric. A homotopy can be constructed by deforming $g$ to the Euclidean metric. In particular, the map $N\colon\Sigma_{\lambda}\to S^{n-1}$ has degree $1$.
Proposition 5.2 (cf. [Brendle]*Proposition 3.9).
@@
We define a function $W_{\lambda}\colon\Sigma_{\lambda}\to\mathbb{R}$ by $W_{\lambda}=V_{\lambda}-|\mathrm{tr}(q)-q(\nu,\nu)|$.
@@
Suppose that the Matching Angle Hypothesis is satisfied. Then
$\sup_{\Sigma_{\lambda}}\max\{-W_{\lambda},0\}\leq o(\lambda)\quad\text{ as }\lambda\to\infty.$
@@
It follows from [Brendle]*Proposition 3.10 that $\sup_{\Sigma_{\lambda}}\max\{-V_{\lambda},0\}\leq o(\lambda)$ as $\lambda\to\infty$. Since $W_{\lambda}-V_{\lambda}$ is uniformly bounded, the assertion follows. ∎
In the remaining part of this section, we will estimate the $L^{\sigma}$-norm $\max\{-W_{\lambda},0\}$ on $\Sigma_{\lambda}\cap B_{r}(p)$, where $\sigma\in[1,\tfrac{3}{2})$ is a fixed exponent and $B_{r}(p)$ denotes a Euclidean ball of radius $r$.
@@
Consider three pairwise distinct elements $i_{1},i_{2},i_{3}\in I$. We denote by $G_{\lambda}^{(i_{1},i_{2},i_{3})}$ the set of all points $x\in\Sigma_{\lambda}$ with the property that $u_{i_{1}}(x)\geq u_{i_{2}}(x)\geq u_{i_{3}}(x)$ and $u_{i_{3}}(x)\geq u_{i}(x)$ for each $i\in I\setminus\{i_{1},i_{2},i_{3}\}$.
@@
For each $i\in I$, we assume that $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$, where $H$ denotes the mean curvature of the hypersurface $\{u_{i}=0\}$ with respect to the metric $g$. Let us fix an exponent $\sigma\in[1,\tfrac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$. If $\lambda r$ is sufficiently large, then
@@
We adapt the proof of [Brendle]*Lemma 3.13. Let us consider an arbitrary point
$x\in G_{\lambda}^{(i_{1},i_{2},i_{3})}\quad\text{ with }u_{i_{2}}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}.$
By definition of $G_{\lambda}^{(i_{1},i_{2},i_{3})}$, it follows that $u_{i}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$ for all $i\in I\setminus\{i_{1}\}$. Using the identity
$\sum_{i\in I}e^{\lambda u_{i}(x)}=1,$
we obtain $e^{\lambda u_{i_{1}}(x)}\geq 1-C\,e^{-(\lambda r)^{\frac{1}{8}}}$, hence $u_{i_{1}}(x)\geq-C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$. Arguing as in [Brendle], we obtain
@@
at the point $x$. Since $u_{i_{1}}(x)\geq-C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$ and $u_{i}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$ for all $i\in I\setminus\{i_{1}\}$, we can find a point $y\in\Omega$ such that
$u_{i_{1}}(y)=0\quad\text{ and }\quad d(x,y)\leq C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}.$
By assumption, $H-|\mathrm{tr}(q)-q(\nu_{i_{1}},\nu_{i_{1}})|\geq 0$ at the point $y$, where $H$ denotes the mean curvature of the hypersurface $\{u_{i_{1}}=0\}$ with respect to the metric $g$. This implies
@@
for each point
$x\in G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}.$
On the other hand, the intersection $\Sigma_{\lambda}\cap B_{r}(p)$ has area at most $Cr^{n-1}$ (see [Brendle]*Lemma 3.11). Consequently,
@@
Assume that the Matching Angle Hypothesis holds. Let us fix an exponent $\sigma\in[1,\tfrac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$. If $\lambda r$ is sufficiently large, then
$\bigg(r^{\sigma+1-n}$ $\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\geq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}\cap\{u_{i_{3}}\leq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}\leq C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}$
@@
The corresponding estimate for $V_{\lambda}$ follows from [Brendle]*Lemma 3.14. Moreover, it is shown in [Brendle] that the set $G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\geq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}\cap\{u_{i_{3}}\leq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}$ has area at most $C\,(\lambda r)^{-\frac{7}{8}}\,r^{n-1}$. Since $|W_{\lambda}-V_{\lambda}|\leq C\leq C\lambda^{\frac{1}{8}}r^{-\frac{7}{8}}$, the assertion follows. ∎
@@
Let us fix an exponent $\sigma\in[1,\tfrac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$. If $\lambda r$ is sufficiently large, then
@@
The corresponding estimate for $V_{\lambda}$ follows from [Brendle]*Lemma 3.15. Moreover, it is shown in [Brendle] that the set $G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{3}}\geq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}\cap B_{r}(p)$ has area at most $C\,(\lambda r)^{-\frac{3}{2}}\,r^{n-1}$. Since $|W_{\lambda}-V_{\lambda}|\leq C\leq C\lambda$, the assertion follows. ∎
@@
For each $i\in I$, we assume that $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$, where $H$ denotes the mean curvature of the hypersurface $\{u_{i}=0\}$ with respect to the metric $g$. Moreover, we assume that the Matching Angle Hypothesis is satisfied. Let us fix an exponent $\sigma\in[1,\tfrac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$. If $\lambda r$ is sufficiently large, then
$\bigg(r^{\sigma+1-n}$ $\int_{\Sigma_{\lambda}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}+C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}+C\,(\lambda r)^{1-\frac{3}{2\sigma}}.$
@@
$\bigg(r^{\sigma+1-n}$ $\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}+C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}+C\,(\lambda r)^{1-\frac{3}{2\sigma}}$
@@
For each $i\in I$, we assume that $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$, where $H$ denotes the mean curvature of the hypersurface $\{u_{i}=0\}$ with respect to the metric $g$. Moreover, we assume that the Matching Angle Hypothesis is satisfied. Let us fix an exponent $\sigma\in[1,\tfrac{3}{2})$. Then
$\sup_{p\in\mathbb{R}^{n}}\sup_{r\leq 1}\bigg(r^{\sigma+1-n}\int_{\Sigma_{\lambda}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}\to 0\quad\text{ as }\lambda\to\infty.$
@@
$(\delta_{l}\lambda_{l})^{-1}\sup_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}\to 0\quad\text{ as }l\to\infty.$
This implies